Project

General

Profile

NDOS Commissioning Meeting 2010/12/20

Attendance: Jon, Holger, Ryan, Sue, Rick K., Ron, Neil, Greg, Mark Messier, Rick T., Peter, Art, Bob, Chad, Kurt, Denis, Mat, Gineidy, Qiming, Leon

Agenda:
  1. DAQ/Electronics Update & Plans
  2. DCS Status & Plans
  3. Shifters' report
  4. AOB

DAQ:

  • Peter:
    • Qiming made some changes to the DCM timing code that should make things more table, in particular this should resolve the FEB-out-of-synch problem we've been having at the start of some runs.
    • A "checkpoint" release of the DAQ software has been implemented, and we ran with that over the past weekend. The current checkpoint release was the development tag of Dec. 16.
    • MiniBoone triggers are still not implemented, but hopefully we'll get to that today.
  • Ron:
    • Would like to start testing the timing system commands, which will involve interrupting the TDU/spillserver. Ron and company will call the Control Room if a reset of the TDU is needed.
    • Last week some tests were done:
      • data drain for high rates on DCM
      • dcm_get_status, code will be put into general DCMProgUtils code.
      • saw a modest use of DCM CPU, even when just logging in
      • might need to put ?? in to the real-time scheduler
  • Sue:
    • Posted a user's guide on the NOvA DAQ Monitor, doc-dB 5474
    • Would like to work on sending Run Control error codes
  • Mat:
    • Last week Denis and Mat investigated most of the channels that have been masked-off.
      • Some channels/FEBs were recovered; these had loose connections or bad cables.
      • Two FEBs were found to be dead and were swapped out.
      • Looked at some APDs that had large baseline noise. Using an electrometer, they found a ~2 minute oscillation that is most likely tied to the thermal cycle of the chiller.
      • 1 APD was found to be loose; this was removed.
      • Seven (7) other APDs that had bad behavior on ~same pixels were removed. Upon a visual inspection, some oil was found in the o-ring and dots and blemishes are seen on the face of the APD. Photos are in the CRL.
      • We have enough (dry) parts to replace 5 APDs, however these will not be replaced until we have a better understanding of what happened and Leon gives the green-light.
      • The bad APDs were found on both the top of diblock 1 and on the side of diblock 2.
  • Rick T:
    • All of the replacement power cables are now at the test-stand.
    • Andrew needs ~1/2 day of work to swap out the cables.
    • We have tentatively planned for this to be done Wed. 9 am
  • Andrew will be given access to the DAQ at 2 pm today to re-do some DSO tests.

DCS:

  • Chad: the rack-protection wiring and read-out is now complete. Rack 2 is reading out racks 1 & 3, rack 1 is reading out rack 2. Plans to work on getting data written to the dB
  • DCM status readback: still cannot read the status bits without interrupting data flow and crashing the DAQ. Work-in-progress.

Shifters' report: nothing specific, the weekend shifts went relatively smoothly (considering)

AOB:

  • There will be no Commissioning meeting this Thursday.
  • Merry Christmas!